# E-Trojans: Ransomware, Tracking, DoS, and Data Leaks on Battery-powered Embedded Systems

### v1

@@ 1 Introduction
Real-world BESs have complex and proprietary internals, including hardware and firmware components. The internal typically consists of a microcontroller, running closed-source firmware, a rechargeable battery, and an electric motor. Moreover, they include a wireless communication interface such as Bluetooth Low Energy (BLE), and internal buses including Universal Asynchronous Receiver-Transmitter (UART) and Inter-Integrated Circuit (I2C). BESs are operated by users and can be controlled using a smart device, e.g., an e-scooter is driven by the user and controlled via a mobile companion app.
@@
We fill this gap by presenting the first security and privacy assessment of the internal attack surface of e-scooters. We focus on e-scooters from Xiaomi and Segway-Ninebot [19], two market leaders. Xiaomi has released seven e-scooters in the last seven years: M365, Pro, Pro 2, 1S, Essential, Mi 3 (ES3), and Mi 4. We target the M365 and Mi 3 (ES3) models, as they cover two e-scooter generations (2016 – 2024) and are used daily by millions of people [60]. These users include people owning private e-scooters and those using rented Xiaomi e-scooters from services like Bird and Lime [6, 70]. Our analysis also covers Mi Home, the Xiaomi e-scooter companion app available for Android [4] and iOS [5]. We uncover four design vulnerabilities by reverse-engineering (RE) the M365 and ES3 internals. The vulnerabilities include unsigned and unencrypted BMS (V1, V2) and unencrypted and unauthenticated UART/I2C communication (V3, V4). Being design flaws, they are effective on any Xiaomi M365 or Mi 3 e-scooters and possibly other Xiaomi e-scooters. We also report novel RE findings, including the e-scooters’ hardware block diagram, firmware protections, and battery management details.
We propose E-Trojans, four novel attacks on BES internals, compromising the safety, security, privacy, and availability of the Xiaomi e-scooter ecosystem. The attacks exploit the vulnerabilities we found to achieve malicious goals, including (remote) code execution on the BMS. Among the attacks, we present the first BES battery ransomware. The attack remotely infects the BMS and progressively and irreversibly destroys the e-scooter’s battery via undervoltage until the victim pays a ransom. We also showcase a novel tracking attack that leverages immutable hardware identifiers within the e-scooter’s internals to track its driver.
@@
We successfully tested the E-Trojans attacks in a realistic but controlled scenario against up-to-date M365 and ES3 e-scooters and report our experimental results. For instance, we were able to degrade the M365 battery’s autonomy by 50% in three hours using the undervoltage battery ransomware PoC and track a user via unique e-scooter fingerprints broadcast over BLE.
To fix the presented vulnerabilities and attacks, we propose four countermeasures (C1–C4). Our fixes add confidentiality and integrity guarantees for the battery controller firmware and protect the internal UART bus against spoofing and DoS. Moreover, they are backward-compliant with the Xiaomi ecosystem, as they reuse crypto primitives already supported, and low-cost because they add minimal overhead.
@@
Disclosure, availability, and ethics. In November 2023, we responsibly disclosed our findings and PoCs to Xiaomi [75]. Xiaomi acknowledged the report and closed it as informative in March 2024. In April 2024, we contacted ST Microelectronics (ST) and Texas Instruments (TI), the manufacturers of the battery management chips we attacked. ST responded that our attacks do not raise any security concerns as they are physical. We disagree with their assessment as we conduct our attacks in wireless proximity or even remotely. TI has not responded. We provide our toolkit and attack demos at https://github.com/Skiti/Etrojans. We ethically conducted our experiments: we tested our e-scooters and smartphones in a laboratory, and minimized safety risks during the undervoltage and overvoltage experiments.
@@ 2.2 Threat Model
System Model. Our system model follows the block diagram in Figure 1. It includes a user who owns a Xiaomi e-scooter (BES) and a smartphone running Mi Home for Android or iOS. They paired Mi Home with the e-scooter and locked it with a password. The user runs the latest version of the BTS, DRV, BCTRL firmware, and Mi Home and charges the e-scooter with the original Xiaomi charger.
Attacker Models. We focus on protocol-level attacks and consider two related attacker models: (i) proximity-based attacker who interacts with the e-scooter wirelessly over BLE (e.g., using a laptop in the BLE range with the e-scooter); (ii) remote attacker who installs a rogue malicious app on the victim’s smartphone and exploits the e-scooter remotely via said app. Attacks at the implementation-level, including side channel and fault injection, are out of scope.
Attacker Goals. The attacker wants to violate the safety, security, privacy, and availability of the e-scooter by exploiting V1, V2, V3, and V4 (i.e., architectural vulnerabilities we found in the internals). For instance, the adversary could try to extort money from the e-scooter driver via ransomware or follow their movements via a tracking attack. These two goals are novel in the domain of e-scooters, leading to large-scale and critical impacts as they target the entire Xiaomi e-scooter ecosystem, affecting millions of users and devices.
@@ 3 Xiaomi E-Scooters RE and Vulnerabilities
We present prior and new details about Xiaomi e-scooter internals found via RE. Then, we describe the four vulnerabilities we uncovered.
@@ 3.1 Available RE Information
E-scooter block diagram. As shown in Figure 1, a Xiaomi e-scooter is composed of three undocumented and closed-source systems: Bluetooth Low Energy (BTS), Driving (DRV), and Battery Management (BMS). BTS provides low-power and reliable wireless communication, DRV controls the electric motor, and BMS manages the battery. Each system has a dedicated System-On-Chip (SoC) running closed-source firmware. The systems communicate using a proprietary protocol over a Universal Asynchronous Receiver-Transmitter (UART) bus. UART provides a serial, asynchronous, and full-duplex digital communication channel with Tx and Rx wires.
Mi Home app. The user interacts with a Xiaomi e-scooter via the Mi Home application for Android [4] or iOS [5]. The e-scooter and the app communicate over BLE using a custom application-layer protocol studied in [12]. They discover each other using BLE scanning and advertising. To connect and communicate with the e-scooter, the user logs in with their Xiaomi account on Mi Home and pair the app and the e-scooter.
Then, the paired devices connect whenever they are in the BLE range without requiring user interaction. Mi Home offers valuable features for controlling the e-scooter. For example, it allows to set a password to lock the e-scooter, perform over-the-air (OTA) firmware updates, and monitor the e-scooter’s battery status.
@@ 3.2 New RE Information
Starting from the details presented in Section 3.1, we RE the internals of the M365 [52] and Mi 3 (ES3) [53] Xiaomi e-scooters. Next, we summarize our findings.
@@
The BCTRL is an STM8L151K6 chip [64] from ST, with OTA firmware updates [63]. We identified the VDD, SWIM, GND, and RST pins used for powering, programming, and resetting the BCTRL. We soldered them to an ST-Link v2 programmer for firmware extraction and debugging, and found no hardware debug (ST-Link) protection on the BCTRL. The BMON is a BQ76930 chip [25] from TI, configurable by the BCTRL. Unlike the BCTRL, the BMON does not support firmware updates or hardware debugging [26].
The BCTRL and BMON communicate using a proprietary protocol over I2C, a synchronous serial digital protocol running on two wires (SDA and SCL). The BCTRL (I2C master) controls the BMON (I2C slave). We found that the proprietary I2C protocol lacks encryption and authentication.
@@
Firmware signing and encryption. We gathered BTS, DRV, and BCTRL firmware binaries from the Mi Home app cache, Xiaomi backend, and public repositories. We also connected to the BCTRL through the SWIM interface, which is used for programming and debugging to dump its current firmware. Our firmware collection ranges from the M365 firmware released in 2019 to the most recent firmware for the ES3 released in 2024. We inspected them with reverse-engineering tools like Ghidra and tested them with techniques like entropy analysis. The M365 firmware was unencrypted and unsigned. The ES3 BTS firmware was signed with ECDSA since v1.5.2, the DRV firmware was signed with ECDSA and encrypted with TEA since v0.1.7, and the BCTRL firmware was unprotected.
@@
Table 1: Battery and BMS details. OVT: overvoltage threshold, UVT: undervoltage threshold, LBD: load balancing delta. Prefix d means dangerous, and prefix c means critical.
M365 ES3
Battery
Manuf. Ninebot Ninebot
Type Li-Ion 18650 Li-Ion 18650
Chem. 10INR19/66-3 10INR19/66-3
Cells 30 (10x3) 30 (10x3)
Capacity 7.80Ah/300Wh 7.65Ah/275Wh
Voltage 42V to 36V 42V to 36V
BMS
BCTRL STM8L151K6 STM8L151K6
BMON BQ76930 BQ76930
dOVT 4200mV 4200mV
cOVT 4700mV 4700mV
dUVT 2750mV 2750mV
cUVT 1580mV 1580mV
dLBD 30mV 30mV
cLBD 800mV 800mV
BMS. We RE the M365 and ES3 BMS by studying the unencrypted BCTRL firmware and the BMON datasheet. The BMON periodically reads the battery’s status, like its voltage and temperature, and relays it to the BCTRL. The BCTRL manages the battery based on the data received from the BMON and initializes the configuration parameters of the BMON, including the undervoltage and overvoltage battery protection thresholds.
The BMS undervoltage threshold is set between 1580mV and 2750mV, and the overvoltage threshold is between 4200mV and 4700mV. We define these intervals as critical and dangerous undervoltage (cUVT = 1580mV, dUVT = 2750mV) and overvoltage (dOVT = 4200mV, cOVT = 4700mV) intervals. If a battery cell voltage exceeds the established threshold, the BCTRL raises an internal error and powers OFF the BTS and DRV, shutting down the e-scooter.
The BMS has a proprietary load-balancing mechanism that is essential for the battery’s health. Load balancing transfers voltage among battery cells, aiming for a state where all cells possess the same voltage. Imbalanced cells can reach a state of undervoltage or overvoltage and get severely and irreversibly damaged and even cause fire hazards and explosions [76]. Load balancing occurs when the e-scooter is charging or in sleep mode (the user powers OFF the e-scooter). According to our RE findings, imbalanced cells exceed the critical load balancing delta (cLBD = 800mV) and cause the BMS to raise an error and power OFF the BTS and DRV.
@@
1:initBCTRL(serial, fwver, dLBD, cLBD)
@@
19:         if battcells.deltaVolt >=>=> = dLBD then
@@
BCTRL firmware. We RE the BCTRL firmware initialization and the main loop by decompiling its stripped binary and debugging its execution with an ST-Link device connected to the SWIM interface, running GDB. As shown in Algorithm 1, during initialization, the BCTRL 1) loads its configuration parameters, including serial number, firmware version, and load balancing deltas; and 2) configures the BMON over I2C by writing into specific BMON registers, including setting the undervoltage and overvoltage thresholds.
Then, the firmware enters a main loop where it: 5) reads the battery cells’ voltage and current values from special BMON read-only registers; 6) computes the current battery level (percentage); 7) checks voltage levels against cLBD, 9) enables load balancing if needed, and 11) checks undervoltage against cUVT and 14) overvoltage against cOVT.
If a critical undervoltage, overvoltage, or balancing check fails, it sends error messages that power OFF the BTS and DRV, and 18) enters sleep mode and performs load balancing. During the main loop, the BCTRL 23) reads from and 24) writes to the UART bus to communicate with the BTS and DRV.
@@ 3.3 New Vulnerabilities
During our RE experiments on the M365 and ES3 e-scoters, we discovered four vulnerabilities:
@@ 1. V1:
    BCTRL firmware is unencrypted An attacker can retrieve the (stripped) BCTRL firmware in plaintext (e.g., during a firmware update) and RE it.
@@ 2. V2:
    BCTRL firmware is unsigned An attacker can modify the firmware of the BCTRL, including its safety thresholds and how it initializes the BMON, and flash the modified BCTRL on the victim’s e-scooter.
@@ 3. V3:
    UART bus lacks integrity protection, encryption, and authentication An attacker with access to the UART bus (e.g., via a malicious BCTRL firmware) can eavesdrop on all UART messages, replay or forge them, and impersonate or MitM the DRV and BTS.
@@ 4. V4:
    UART bus lacks protection against DoS An attacker accessing the UART bus can DoS internal (BMS, BTS, and DRV) and external (Mi Home and battery charger) components.
These issues belong to known vulnerabilities classes, but are novel in the context of e-scooters. They stem from design flaws in Xiaomi proprietary protocols, independent of specific hardware or implementation details. As a result, our four vulnerabilities impact the entire Xiaomi ecosystem, potentially affecting other Xiaomi devices using a similar BCTRL and UART bus.

### v2

@@ 1 Introduction
Real-world BESs have complex and proprietary internals, including hardware and firmware components. The internal typically consists of a microcontroller, running closed-source firmware, a rechargeable battery, and an electric motor. Moreover, they include a wireless communication interface such as Bluetooth Low Energy (BLE), and internal buses including Universal Asynchronous Receiver-Transmitter (UART) and Inter-Integrated Circuit (I2C). BESs are operated by users and can be controlled using a smart device, e.g., an e-scooter is ridden by the user and controlled via a mobile companion app.
@@
We fill this gap by presenting the first security and privacy assessment of the internal attack surface of e-scooters. We focus on e-scooters from Xiaomi and Segway-Ninebot [21], two market leaders. Xiaomi has released seven e-scooters in the last seven years: M365, Pro, Pro 2, 1S, Essential, Mi 3 (ES3), and Mi 4. We target the M365 and Mi 3 (ES3) models, as they cover two e-scooter generations (2016 – 2024) and are used daily by millions of people [66]. These users include people owning private e-scooters and those using rented Xiaomi e-scooters from services like Bird and Lime [6, 77]. Our analysis also covers Mi Home, the Xiaomi e-scooter companion app available for Android [4] and iOS [5].
We uncover four design vulnerabilities by reverse-engineering (RE) the M365 and ES3 internals. The vulnerabilities include unsigned and unencrypted BMS (V1, V2) and unencrypted and unauthenticated UART/I2C communication (V3, V4). Being design flaws, they are effective on any Xiaomi M365 or Mi 3 e-scooters and possibly other Xiaomi e-scooters. We also report novel RE findings, including the e-scooters’ hardware block diagram, firmware protections, and battery management details.
We propose E-Trojans, four novel attacks on BES internals, compromising the safety, security, privacy, and availability of the Xiaomi e-scooter ecosystem. The attacks exploit the vulnerabilities we found to achieve malicious goals, including (remote) code execution on the BMS. Among the attacks, we present the first BES battery ransomware. The attack remotely infects the BMS and progressively and irreversibly destroys the e-scooter’s battery via undervoltage until the victim pays a ransom. We also showcase a novel tracking attack that leverages immutable hardware identifiers within the e-scooter’s internals to track its rider.
@@
We successfully tested the E-Trojans attacks in a realistic but controlled scenario against up-to-date M365 and ES3 e-scooters and report our experimental results. For instance, we were able to degrade the M365 battery’s autonomy by 50% in three hours using the undervoltage battery ransomware PoC and to track a user via unique e-scooter fingerprints broadcasted over BLE.
To fix the presented vulnerabilities and attacks, we propose four countermeasures (C1–C4). Our fixes add confidentiality and integrity guarantees for the battery controller firmware and protect the internal UART bus against spoofing and DoS. Moreover, they are backward-compatible with the Xiaomi ecosystem, as they reuse cryptographic primitives already supported, and low-cost because they add minimal overhead.
@@
Responsible Disclosure. We followed standard practices for responsible disclosure by contacting Xiaomi and the affected vendors. We notified Xiaomi in November 2023 via their official HackerOne bug bounty program [82]. We uploaded and shared with them video demonstrations, PoCs, and instructions on how to reproduce the attacks. Xiaomi acknowledged our multiple reports, updated its CVSS severity score to Medium (5.8), and closed it as informative in March 2024. We disagreed with Xiaomi’s assessment and, to have our findings addressed, in April 2024 we contacted via email the manufacturers of the BMS chips we targeted: STMicroelectronics (ST) and Texas Instruments (TI). ST responded that our attacks do not raise any security concerns as they are physical. We contest their assessment as we conduct our attacks in wireless proximity or even remotely. TI did not respond.
Availability. We provide our toolkit and attack demos at redacted. For safety reasons, we redacted the code responsible for generating malicious BMS firmware, which took months of RE to develop. We also release a patched BCTRL firmware that blocks BMS firmware updates, preventing our attacks. The private repository will be open-sourced in the future.
Ethics. We conducted our experiments ethically: we tested our e-scooters and smartphones in a laboratory, and minimized safety risks during the undervoltage and overvoltage experiments.
@@ 2.2 Threat Model
System Model. Our system model follows the block diagram in Figure 1. It includes a user who owns a Xiaomi e-scooter (BES) and a smartphone running Mi Home for Android or iOS. They paired Mi Home with the e-scooter and locked it with a password. The user runs the latest version of the BTS, DRV, BCTRL firmware, and Mi Home, and charges the e-scooter with the original Xiaomi charger. The Xiaomi ecosystem includes e-scooters owned by a user or rented from a fleet. Our model includes both personal and fleet e-scooters and we focus on their Xiaomi-specific features. We do not study features added by a fleet manager on top of the Xiaomi e-scooter, like a cellular module with a SIM card.
Attacker Models. We focus on protocol-level attacks and consider two related attacker models: (i) proximity-based attacker who interacts with the e-scooter wirelessly over BLE (e.g., using a laptop in the BLE range with the e-scooter); and a (ii) remote attacker who can connect to the e-scooter over the Internet through malware installed on a victim’s device (e.g., using a malicious Android app). Attacks at the implementation-level, including side channel and fault injection, are out of scope.
Our proximity-based and remote attacker models correspond to several real-world attack scenarios. For example, the proximity-based attacker can place a (cheap) BLE transmitter in a public location (e.g., a train station or a parking spot) to target e-scooters in transit, or attach it beneath an e-scooter, to attack nearby e-scooters as it moves. Similarly, the remote attacker can install a malicious app on a victim’s smartphone to attack their e-scooter, potentially performing timing attacks, such as triggering errors while the user is riding it.
Attacker Goals. The attacker wants to violate the safety, security, privacy, and availability of the e-scooter by exploiting V1, V2, V3, and V4 (i.e., architectural vulnerabilities we found in the internals). For instance, the adversary could try to extort money from the e-scooter’s rider via ransomware or follow their movements via a tracking attack. These two goals are novel in the domain of e-scooters, leading to large-scale and critical impacts as they target the entire Xiaomi e-scooter ecosystem, affecting millions of users and devices.
@@ 3 Xiaomi E-Scooters RE and Vulnerabilities
We present prior and new details about Xiaomi e-scooter internals we found via RE. Then, we describe the four vulnerabilities we uncovered.
@@ 3.1 Available RE Information
E-Scooter Block Diagram. As shown in Figure 1, a Xiaomi e-scooter is composed of three undocumented and closed-source systems: Bluetooth Low Energy (BTS), Driving (DRV), and Battery Management (BMS). BTS provides low-power and reliable wireless communication, DRV controls the electric motor, and BMS manages the battery. Each system has a dedicated System-On-Chip (SoC) running closed-source firmware. The systems communicate using a proprietary protocol over a Universal Asynchronous Receiver-Transmitter (UART) bus. UART provides a serial, asynchronous, and full-duplex digital communication channel with TX/RX wires.
Mi Home App. The user interacts with a Xiaomi e-scooter via the Mi Home application for Android [4] or iOS [5]. The e-scooter and the app communicate over BLE using a custom application-layer protocol studied in [12]. They discover each other using BLE scanning and advertising.
To connect and communicate with the e-scooter, the user logs in with their Xiaomi account on Mi Home and pairs the app and the e-scooter. Then, the paired devices connect whenever they are in the BLE range without requiring user interaction.
Mi Home offers valuable features for controlling the e-scooter. For example, it allows to set a password to lock the e-scooter, perform over-the-air (OTA) firmware updates, and monitor the e-scooter’s battery status.
@@ 3.2 New RE Information
Starting from the details presented in Section 3.1 and known information about the internals of the M365 [58], we RE the Mi 3 (ES3) [59] Xiaomi e-scooter and discuss the parts relevant to our attacks. Next, we summarize our findings.
@@
The BCTRL is an STM8L151K6 chip [70] from ST and supports firmware updates [69]. We identified the VDD, SWIM, GND, and RST pins used for powering, programming, and resetting the BCTRL. We soldered them to an ST-Link v2 programmer for firmware extraction and debugging and found no hardware debug (ST-Link) protection on the BCTRL. The BMON is a BQ76930 chip [30] from TI, configurable by the BCTRL. Unlike the BCTRL, the BMON does not support firmware updates or hardware debugging [31].
The BCTRL and BMON communicate using a proprietary protocol over I2C, a synchronous serial digital protocol running on two wires (SDA and SCL). The BCTRL (I2C master) controls the BMON (I2C slave). The I2C protocol lacks encryption and authentication.
@@
Firmware Signing and Encryption. We gathered BTS, DRV, and BCTRL firmware binaries from the Mi Home app cache, the Xiaomi backend, and a public repository. We also connected to the BCTRL through the SWIM interface, which is used for programming and debugging, to dump its current firmware. Our firmware collection ranges from the M365 firmware released in 2019 to the most recent firmware for the ES3 released in 2024. We statically analyze them with Ghidra to study their encryption and signature verification. We discover that the BTS, DRV, and BCTRL firmware from M365 are unencrypted and unsigned, regardless of their version. Conversely, in the ES3, the BTS firmware is signed with ECDSA since v1.5.2, the DRV firmware is signed with ECDSA and encrypted with TEA since v0.1.7, and the BCTRL firmware is unprotected.
@@
BMS. We RE the e-scooter’s BMS by focusing on the unencrypted BCTRL firmware and the BMON datasheet. The BCTRL and BMON work together to ensure safe battery management. Our analysis focuses on three unsafe conditions that impact battery health and safety: undervoltage, which can lead to short circuits and permanent battery damage; overvoltage, which increases the risk of overheating, fire, and thermal runaway; and voltage imbalance between battery cells, which contributes to both undervoltage and overvoltage, accelerating battery degradation.
The BMON measures battery parameters, such as voltage and temperature, reports fault conditions to the BCTRL, and stores thresholds that activate hardware protections. The undervoltage threshold can be set between 1580mV and 2750mV, while the overvoltage threshold can be set between 4200mV and 4700mV. We consider them critical and dangerous undervoltage thresholds, (cUVT = 1580mV and dUVT = 2750mV) and critical and dangerous overvoltage thresholds, (cOVT = 4700mV and dOVT = 4200mV).
The BCTRL is responsible for the battery’s health, processing data from the BMON and determining whether safety mechanisms should be triggered. It has write access to the BMON hardware protection registers, allowing it to modify safety thresholds and disable safety mechanisms. The BCTRL also manages load balancing, ensuring voltage is evenly distributed among individual battery cells by defining the critical load balancing delta (cLBD = 800mV), which sets the maximum allowable voltage difference between any two cells. Additionally, the BCTRL handles communication with other e-scooter components, such as BTS and the DRV, and powers them off when the battery enters an unsafe state.
@@
1:initBCTRL(serial, fwver, cLBD)
@@
19:         if battcells.deltaVolt >=>=> = cLBD then
@@
BCTRL Firmware. We RE the BCTRL firmware initialization and main loop by decompiling its stripped binary and debugging it with an ST-Link device connected to the SWIM interface. As shown in Algorithm 1, during initialization, the BCTRL 1) loads its configuration parameters, including serial number, firmware version, and load balancing deltas; and 2) configures the BMON over I2C by writing into related registers, including setting the undervoltage and overvoltage thresholds.
Then, the BCTRL firmware enters a main loop where it: 5) reads the battery cells’ voltage and current values from special BMON read-only registers; 6) computes the current battery level (percentage); 7) checks voltage levels against cLBD, 9) enables load balancing if needed, and 11) checks undervoltage against cUVT and 14) overvoltage against cOVT.
If a critical undervoltage, overvoltage, or balancing check fails, the BCTRL sends error messages that power off the BTS and the DRV, 18) enters sleep mode, 20) and performs load balancing. During the main loop, the BCTRL 23) reads from and 24) writes to the UART bus to communicate with the BTS and the DRV.
@@ 3.3 New Vulnerabilities
During our RE experiments on the M365 and ES3 e-scooters, we discovered four vulnerabilities:
@@ 1. V1:
    BCTRL firmware is unencrypted. An attacker can retrieve the (stripped) BCTRL firmware in plaintext (e.g., during a firmware update) and RE it.
@@ 2. V2:
    BCTRL firmware is unsigned. An attacker can modify the firmware of the BCTRL, including its safety thresholds and how it initializes the BMON, and flash the modified BCTRL on the victim’s e-scooter.
@@ 3. V3:
    UART bus lacks integrity protection, encryption, and authentication. An attacker with access to the UART bus (e.g., via a malicious BCTRL firmware) can eavesdrop on all UART messages, replay or forge them, and impersonate or MitM the DRV and BTS.
@@ 4. V4:
    UART bus lacks protection against DoS. An attacker accessing the UART bus can DoS internal (BMS, BTS, and DRV) and external (Mi Home and battery charger) components.
These issues belong to known vulnerability classes but are novel in the context of e-scooters. They stem from design flaws in Xiaomi proprietary protocols, independent of specific hardware or implementation details. As a result, our four vulnerabilities impact the entire Xiaomi ecosystem, potentially affecting other Xiaomi devices using a similar BCTRL and UART bus.